Santa Klaus Toys just paid a dividend of $3.00 per share. The required return is 11.7 percent and the perpetual dividend growth rate is 3.9 percent. What price should this stock sell for five years from today?
Answer #1

D0=3

Price of the stock 5 years from today P5=D6/(r-g)=D0*(1+g)^6/(r-g)=3*(1+3.9%)^6/(11.7%-3.9%)=48.3860241